# Multi-Path and Multi-Particle Tests of the Dimensionality of Quantum Mechanics

### v1

@@ Conclusion.
In summary, we derived in a most general way the Peres test via a matrix formalism, also allowing for a geometrical interpretation of the test. We then introduced generalized Peres tests that exploit multi-path and multi-particle interference, revealing a direct relation to the dimension of the number system of quantum mechanics. Future theoretical works have to explore, how multifaceted interference of mutually incoherent particles modifies these tests and its sensitivity to higher dimensions; and future experiments have to show whether complex numbers are not only necessary but also sufficient for quantum mechanics.
@@ Acknowledgments.
E.İ.S. thanks Miguel Navascués for fruitful discussions and gratefully acknowledges financial support by the International Max Planck Research School: Physics of Light (IMPRS-PL). M.-O.P. and J.v.Z. acknowledge funding by the Erlangen Graduate School in Advanced Optical Technologies (SAOT) by the Bavarian State Ministry for Science and Art. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”).

### v2

@@ Conclusion.
In summary, we derived in a most general way the Peres test via a matrix formalism, what allows in particular for a geometrical interpretation of the test. We further introduced generalized Peres tests that exploit multi-path and multi-particle interference, revealing a direct relation to the dimension of the number system of quantum mechanics. Future theoretical works have to explore, how multifaceted interference in general, e.g., using mutually incoherent particles modifies these tests and its sensitivity to higher dimensions; furthermore, future experiments have to show whether complex numbers are not only necessary but also sufficient for quantum mechanics.
@@ Acknowledgments.
E.İ.S. thanks Miguel Navascués for fruitful discussions and gratefully acknowledges financial support by the International Max Planck Research School: Physics of Light (IMPRS-PL). J.v.Z. and M.-O.P. acknowledge funding by the Erlangen Graduate School in Advanced Optical Technologies (SAOT) by the Bavarian State Ministry for Science and Art. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”).